Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Gmat prep Question

Expert replies
by rakaisraka » Wed Aug 26, 2015 11:53 am
Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelop with its correct addresses. If the 4 leters are to be put into the 4 envelops at random , what is the probability that only 1 letter will be put into the envelop with its correct address.
a) 1/24
b) 1/8
c) 1/4
d) 1/3
e) 3/8
Join the discussion
Source: — Problem Solving |

by kvamsy » Wed Aug 26, 2015 12:32 pm
I think, Total ways of selecting the envolpes and letters are 4*4 ways.

Not getting the right letter into the envolope is 4+3+2+1 ways into fowur letters.

Total probability of not getting the letter into rt envolope is 10/16 ways.

Probability of getting rt letter into the rt envolope is 1 - 10/16 = 6/16, 3/8

Correct me if i am wrong.
Join the discussion

by Max@Math Revolution » Thu Aug 27, 2015 3:18 am
Forget conventional ways of solving math questions. In PS, IVY approach is the easiest and quickest way to find the answer.

Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelop with its correct addresses. If the 4 leters are to be put into the 4 envelops at random , what is the probability that only 1 letter will be put into the envelop with its correct address.
a) 1/24
b) 1/8
c) 1/4
d) 1/3
e) 3/8

==> probability = number of specific event / total number of cases

Total number of cases = letters(A,B,C,D), envelops(a,b,c,d) thus becomes 4*3*2*1.(A has 1 to choose from a,b,c,d, making it 4, while B has 3, C has 2 and D have 1)

The specific event has the cases (A,a) with (B,d) for (A,a), (C,b), (D,c) or (B,c), (C,d), (D,b), thus there is only 2 case (Since there should be only one letter in the envelope with the exact address)

Since the case applies to (B,b), (C,c), (D,d) as well, the number of cases that only 1 letter will be in the envelope with the exact address =2*4=8.

Thus, probability = 8/4*3*2*1=1/3. Therefore the answer is D.


www.mathrevolution.com
l The one-and-only World’s First Variable Approach for DS and IVY Approach for PS that allow anyone to easily solve GMAT math questions.

l The easy-to-use solutions. Math skills are totally irrelevant. Forget conventional ways of solving math questions.

l The most effective time management for GMAT math to date allowing you to solve 37 questions with 10 minutes to spare

l Hitting a score of 45 is very easy and points and 49-51 is also doable.

l Unlimited Access to over 120 free video lessons at https://www.mathrevolution.com/gmat/lesson

Our advertising video at https://www.youtube.com/watch?v=R_Fki3_2vO8
Join the discussion

by GMATGuruNY » Thu Aug 27, 2015 3:37 am
rakaisraka wrote:Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelop with its correct addresses. If the 4 leters are to be put into the 4 envelops at random , what is the probability that only 1 letter will be put into the envelope with its correct address.
a) 1/24
b) 1/8
c) 1/4
d) 1/3
e) 3/8
Let the 4 letters be A, B, C and D.
Total ways to arrange the 4 letters = 4! = 24.
Let the correct ordering of the 4 letters be ABCD.

Write out the ways that ONLY A can be put in the correct position:
ACDB
ADBC
Total ways = 2.

Using the same reasoning, there will be 2 ways that ONLY B can be put in the correct position, 2 ways that ONLY C can be put in the correct position, and 2 ways that ONLY D can be put in the correct position.
Thus, the total number of ways to put EXACTLY 1 letter in the correct position = 2+2+2+2 = 8.

Thus:
P(exactly 1 letter is put in the correct position) = 8/24 = 1/3.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Brent@GMATPrepNow » Sun Dec 29, 2019 9:53 am
rakaisraka wrote:Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelop with its correct addresses. If the 4 leters are to be put into the 4 envelops at random , what is the probability that only 1 letter will be put into the envelop with its correct address.
a) 1/24
b) 1/8
c) 1/4
d) 1/3
e) 3/8
Let's solve this question using counting methods.

So, P(exactly one letter with correct address) = (number of outcomes in which one letter has correct address)/(total number of outcomes)

Let a, b, c and d represent the letters, and let A, B, C and D represent the corresponding addresses.
So, let's list the letters in terms of the order in which they are delivered to addresses A, B, C, and D.
So, for example, the outcome abcd would represent all letters going to their intended addresses.
Likewise, cabd represent letter d going to its intended address, but the other letters not going to their intended addresses.

-------------------
total number of outcomes
The TOTAL number of outcomes = the number of different ways we can arrange a, b, c, and d

RULE: We can arrange n different objects in n! ways.
So, we can arrange the four letters in 4! ways = 24 ways
-------------------

number of outcomes in which one letter has correct address
Now let's list all possible outcomes in which exactly ONE letter goes to its intended addresses:
- acbd
- adbc
- cbda
- dbac
Aside: At this point you might recognize that there are two outcomes for each arrangement in which one letter goes to its intended address
- dacb
- bdca
- bcad
- cabd
There are 8 such outcomes.
-------------------

So, P(exactly one letter with correct address) = 8/24 = 1/3

Answer: D

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion